OpenStudy (anonymous):

3 Tan[θ] + Cot[θ] == 5 Csc[θ]

OpenStudy (anonymous):

you need to remember\[\tan (\theta)\] is the same as \[\frac{ \sin \theta }{ \cos \theta } \]

OpenStudy (anonymous):

\[\frac{ \sin(\theta) }{ \cos(\theta) }+\frac{ \cos(\theta) }{ \sin(\theta) } \] and you should be able to take it from here possibly. You just combine like terms and remember what csc(theta) is